Along the x-axis, a box is pulled by a string. 30° from the x-axis a second string pulls the box slightly upward. Find the magnitude of both Force vectors.


1
Expert's answer
2021-03-14T19:13:16-0400

Answer

Asume that force "F_1" is upward force and "F_2" is force that working at 30° to the horizontal.

So angle between both force will be

90-30=60°

So magnitude of bet force vector

"F_{net}=\sqrt{F_1^2+F_2^2+2F_1F_2cos60°}\\F_{net}=\sqrt{F_1^2+F_2^2+F_1F_2}"